# A mixed Petrov–Galerkin Cosserat rod finite element formulation

### v2

@@ 1 Introduction
The theoretical foundation of rod mechanics dates back to the early works of [25], who introduced the elastica to describe planar, shear-rigid, and inextensible rods. Later, [39] extended this theory to spatial rods, while [41] generalized it to extensible cases, leading to the term Kirchhoff–Love rod for spatial, nonlinear, shear-rigid structures. Further developments in shear-deformable rod formulations were made by [16], [66], [53], and [62]. Depending on the chosen literature, these rods are referred to as special Cosserat rods [[, see]]Antman1995, Simo–Reissner beams [[, see]]Meier2019, spatial Timoshenko beams [[, see]]Eugster2015, or geometrically exact beams [[, see]]Betsch2002. In this work, we use the term Cosserat rod to describe shear-deformable rods where the centerline and the cross-section orientation are the independent fields of interest. In addition, numerous rod finite element formulations have been proposed in the literature. A comprehensive discussion of their historical evolution and subtle differences is beyond the scope of this work. For an extensive review of both shear-deformable and shear-rigid rod formulations, we refer the interested reader to the survey by [49]. However, to provide context for our approach, we highlight key developments that represent significant milestones in the progression toward the present formulation. In particular, the results of this work build upon and extend the findings of [34] and [24], further contributing to the ongoing development of advanced rod finite element methods.
@@
One of the fundamental challenges in rod formulations is the description of the cross-section orientation when it comes to large rotations. As the cross-section orientation is represented by an orthonormal transformation matrix, which is an element of the special orthogonal group SO(3){SO}(3) and can be parametrized by at least 3 real numbers, some parametrizations suffer from singularities and non-uniqueness. Examples are Euler-angles, Tait–Bryan angles or Rodrigues parameters. To avoid large rotations within one load increment or time step, early formulations employed updated Lagrangian schemes [[, see]]Cardona1988,Ibrahimbegovic1997,Simo1986. While these approaches avoid coming closer to singularities in the parametrization, they require specific update procedures to be handled. Furthermore, they might lead to path-dependent solutions, as noticed by [17]. In contrast, we pursue a total Lagrangian formulation, where all quantities are consistently defined with respect to an inertial basis. For a total Lagrangian formulation, widely known rotation parametrizations used in rod finite element formulations are rotation vectors [13, 34], quaternions [28, 70, 33, 67] and the 9-parameter method [9, 55, 32]. Each parametrization poses additional requirement, such as a complementary update for rotation vectors, a constraint for the length of the quaternion and orthonormality constraints for the 9-parameter method.
@@
Furthermore, an additional challenge arises in the requirement for an objective interpolation of the kinematics. [17] discovered, that the direct interpolation of the nodal (total) rotation vectors leads to a non-objective formulation. Interpolating relative rotation vectors that describe the rotation relative to the first node solves this issue but introduces the additional computational step of evaluating these relative rotation vectors [[, see]]Jelenic1999.
@@
Another key challenge of rod finite element formulations is locking. As discussed in [5], shear and membrane locking in rods can occur if Kirchhoff (shear-rigidity) and inextensibility constraints follow in the limit case of a parameter tending to zero. This appears for instance for very slender rods if the stiffness parameters are computed in the sense of Saint–Venant by using the material’s Young’s and shear moduli, respectively, together with the cross-section geometry. Finite elements that are prone to locking cannot fulfill these constraints exactly over the entire element and introduce parasitic dilatation and shear strains. A common approach to mitigate this issue is reduced integration, which decreases deformation errors but fluctuations in resultant contact forces and moments are still present [[, see for instance]]Meier2015. A reinterpolation of the resultant contact forces and moments is possible to decrease the fluctuations. Other strategies to completely eliminate locking are B-bar methods [[, see for instance]]Greco2017, mixed formulations and three-field formulations based on the Hellinger–Reissner or the Hu–Washizu principle, respectively. Mixed and three-field formulations can be found in plate and shell finite element formulations [[, see]]Atluri1984,Betsch2016, Kirchhoff–Love rods [[, see]]Meier2014,Greco2024, or Cosserat rods [[, see]]Kim1998,Santos2010,Santos2011. In a planar study of arches, [38] use a cubic interpolation of the kinematic fields and quadratic interpolation for the resultant stresses together with the Bubnov–Galerkin approach. In [61, 60], a so-called hybrid-mixed finite element formulation was derived, resulting in different polynomial degrees for the interpolation of orientation, displacement, force and moment. Mixed formulations can also be found in isogeometric analysis of Cosserat rods, see [68, 46], where the resultant contact forces and moments are interpolated with components in the inertial fixed basis and the collocation method is used. Also in rod formulations with enhanced kinematics, where warping or extension of the cross-section is taken into account, mixed formulations are applied, see for example [14, 15].
@@
Mixed and three-field methods do not only solve the problem of locking, they can also significantly improve computational performance. This has been noticed by various authors and with different applications [43, 42, 52, 46, 6]. From a purely numerical perspective for solving systems of nonlinear equations, [1] introduced the so-called lifted Newton method as a strategy to improve the convergence of Newton’s method, when applied to nonlinear root finding problems. Motivated by the multiple shooting method, they highlight, that a better rate of convergence can be achieved, if intermediate variables are introduced as additional degrees of freedom, constrained to their forward computed value. In a mixed formulation, the independent resultant contact forces and moments take the role of these intermediate variables. Since the equivalence of the mixed and displacement-based formulations can be shown under certain circumstances, as demonstrated in [45, 51], the lifted Newton method arises naturally from the mixed formulation.
@@
Another effective way to remove locking is to employ intrinsically locking-free formulations, such as the SE(3){SE}(3) interpolation which has piece-wise constant strains. However, the equations are highly nonlinear, involve trigonometric functions for interpolation, couple position and rotation interpolations, and have computationally expensive evaluations of the force residuals and their derivatives. Additionally, these approaches can impose restrictions on scalability, since a refinement with a higher polynomial degree is not straightforwardly possible. While the authors of this paper used the SE(3){SE}(3) interpolation in a total Lagrangian framework [[, see]]Harsch2023, the interpolation can be done in the SE(3){SE}(3) Lie group setting [[, see]]Sonneville2014, requiring specialized Lie group solvers. [59] also presented a total Lagrangian SE(3){SE}(3) formulation. In contrast to the work of Harsch, where a Petrov–Galerkin projection was applied on the weak form of the Cosserat rod’s equilibrium equations, Santana derived a so-called equilibirum-based formulation, where the resultant contact forces and moments are obtained from the strong form of the rod’s differential equations. Furthermore, strain-based formulations should be mentioned, as they also can represent vanishing strains. They are parametrized directly by the strain measures instead of positions and orientations. A well known approach is the piece-wise constant strain formulation [[, see for instance]]Renda2018, which is widely used in the field of soft robotics. However, a major drawback of strain-based formulations is that positions and orientations can only be obtained by recursive algorithms, leading to asymmetric formulations.
@@
Beyond these numerical challenges, there is also the question of if and how to incorporate constrained rod theories within the general Cosserat rod framework. One approach, as explored in [32], introduces explicit constraints to enforce the Kirchhoff–Love assumptions. Of course, strain-based rod formulations can fulfill the constraints without any further effort, but the previously mentioned drawback remains. The common alternative is to intrinsically satisfy the Kirchhoff–Love assumption by advanced kinematic interpolation itself. Several formulations follow this path, yet they face similar challenges as Cosserat rods do, regarding objectivity, locking, and equation complexity. Moreover, intrinsically Kirchhoff–Love rods are often formulated within an updated Lagrangian framework, inheriting the associated numerical difficulties [[, see for instance]]Meier2016.
@@
The remainder of this paper is organized as follows. In Section˜2, the Cosserat rod theory is briefly recapitulated. It includes the spatially continuous formulations of the internal virtual work and the external virtual work. Two formulations for the internal virtual work are introduced. The first formulation uses only kinematics quantities, while the second is derived using the variational Hellinger–Reissner principle resulting in a formulation that incorporates kinematics and independent resultant contact forces and moments. In Section˜3, the Petrov–Galerkin rod finite element formulation of the Cosserat rod is derived. Therein, the ansatz functions for kinematics, the resultant contact forces and moments, and the corresponding test functions are discretized. The final rod finite element formulation is then obtained by inserting the discretized fields into the continuous formulation from Section˜2, leading to the equations of the generalized force equilibrium. Section˜4 is dedicated to comparing the obtained rod finite element formulations. The comparison includes the displacement-based SE(3){SE}(3) interpolation outlined in [34], and the proposed mixed quaternion formulation of this paper with different polynomial degrees. For completeness, we also consider a mixed SE(3){SE}(3) interpolation and displacement-based quaternion interpolations. Further, we investigate the influence of different spatial integration schemes. Well-studied benchmark experiments are conducted to show the validity and correctness of the proposed formulations. Finally, Section˜5 presents the concluding remarks. In Appendix˜A, the boundary value problem of the Cosserat rod is formulated. Appendix˜B gives further details on the finite element formulation. The detailed derivation of the tangent operator to the quaternion function for non-unit quaternions is given in Appendix˜C. Information on the numerical convergence analyses are given in Appendix˜D.
@@ 2 Cosserat rod theory
As in [62, Equation 2.10], we introduce the diagonal elasticity matrices ${\bm{C}}_{{\bm{\gamma}}}=\operatorname{\operatorname{diag}}(k_{\mathrm{e}},k_{\mathrm{s}_{y}},k_{\mathrm{s}_{z}})$ and ${\bm{C}}_{{\bm{\kappa}}}=\operatorname{\operatorname{diag}}(k_{\mathrm{t}},k_{\mathrm{b}_{y}},k_{\mathrm{b}_{z}})$ with constant coefficients. In the following, the simple quadratic strain energy density
@@
Note that even in the inextensible case, where no strain energy density $W$ with finite stiffness coefficients is available, the complementary strain energy density ${W}^{\star}$ still has finite coefficients. This can be seen as the limit of stiffness components approaching infinity, which results in zeros on the diagonals of ${\bm{C}}_{{\bm{\gamma}}}^{-1}$ and ${\bm{C}}_{{\bm{\kappa}}}^{-1}$. Shear-rigidity can therefore easily be introduced by using ${\bm{C}}_{{\bm{\gamma}}}^{-1}=\operatorname{\operatorname{diag}}(k_{\mathrm{e}}^{-1},0,0)$. Additionally, adding the inextensibility constraint leads to ${\bm{C}}_{{\bm{\gamma}}}^{-1}=\mathbf{0}_{3\times 3}$. The same results were obtained by [23], where still the internal virtual work from ˜4 with the constitutive equations ˜3 was used for the unconstrained deformations.
@@ 3.1 Quaternion interpolation
Its derivation is shown in Appendix˜C. This kinematic interpolation was already proposed by [56], who also provided a proof of objectivity. In contrast to the $\mathbb{R}^{12}$ interpolation introduced by [9] and [55], the quaternion map in ˜13 ensures that ${\bm{A}}_{IB}$ is always an orthonormal transformation matrix.
@@ 3.3 Equations of static equilibrium
are obtained. Note that the number of equations equals in both formulations the number of unknowns, since ${\bm{q}}\in\mathbb{R}^{7N}$ and ${\bm{\lambda}}_{c}\in\mathbb{R}^{6pn_{\mathrm{el}}}$ with the relation $N=pn_{\mathrm{el}}+1$. Following [27], prescribed boundary conditions can be incorporated into the principle of virtual work using perfect bilateral constraints. In both formulations, the nonlinear equilibrium equations can be represented in the form ${\bm{f}}({\bm{x}})=\mathbf{0}_{n}$, where ${\bm{x}}\in\mathbb{R}^{n}$ contains all unknowns, i.e., the generalized coordinates, possibly Lagrange multipliers for bilateral constraints and in the mixed formulation also the generalized forces. Such a nonlinear equation is solved by any root-finding algorithm, e.g., Newton–Raphson, Riks. Note that a system of linear equations with a non-symmetric iteration matrix $\partial{\bm{f}}/\partial{\bm{x}}\in\mathbb{R}^{n\times n}$ must be solved in each iteration. In this work, the static equilibrium will be computed using Newton’s method. As termination criteria we require the Euclidean norm $\sqrt{{\bm{f}}({\bm{x}})^{\!\mathrm{T}}{\bm{f}}({\bm{x}})}$ to be smaller than $\varepsilon\sqrt{n}$, where $\varepsilon$ is the tolerance.
Although we derived only the equations for static equilibrium in this work, for dynamics the equations of motion can also be derived. Since the same generalized coordinates are chosen as in [33] and also the interpolation of the variations is identical, the inertial virtual work and the kinematic differential equation can be derived in the same way, leading to equations of motion.

### v3

@@ 1 Introduction
The theoretical foundation of rod mechanics dates back to the early works of [undefx], who introduced the elastica to describe planar, shear-rigid, and inextensible rods. Later, [undefal] extended this theory to spatial rods, while [undefan] generalized it to extensible cases, leading to the term Kirchhoff–Love rod for spatial, nonlinear, shear-rigid structures. Further developments in shear-deformable rod formulations were made by [undefo], [undefaam], [undefaz], and [undefaai]. Depending on the chosen literature, these rods are referred to as special Cosserat rods [[, see]]Antman1995, Simo–Reissner beams [[, see]]Meier2019, spatial Timoshenko beams [[, see]]Eugster2015, or geometrically exact beams [[, see]]Betsch2002. In this work, we use the term Cosserat rod to describe shear-deformable rods where the centerline and the cross-section orientation are the independent fields of interest. In addition, numerous rod finite element formulations have been proposed in the literature. A comprehensive discussion of their historical evolution and subtle differences is beyond the scope of this work. For an extensive review of both shear-deformable and shear-rigid rod formulations, we refer the interested reader to the survey by [undefav]. However, to provide context for our approach, we highlight key developments that represent significant milestones in the progression toward the present formulation. In particular, the results of this work build upon and extend the findings of [undefag] and [undefw], further contributing to the ongoing development of advanced rod finite element methods.
@@
One of the fundamental challenges in rod formulations is the description of the cross-section orientation when it comes to large rotations. As the cross-section orientation is represented by an orthonormal transformation matrix, which is an element of the special orthogonal group SO(3){SO}(3) and can be parametrized by at least 3 real numbers, some parametrizations suffer from singularities and non-uniqueness. Examples are Euler-angles, Tait–Bryan angles or Rodrigues parameters. To avoid large rotations within one load increment or time step, early formulations employed updated Lagrangian schemes [[, see]]Cardona1988,Ibrahimbegovic1997,Simo1986. While these approaches avoid coming closer to singularities in the parametrization, they require specific update procedures to be handled. Furthermore, they might lead to path-dependent solutions, as noticed by [undefp]. In contrast, we pursue a total Lagrangian formulation, where all quantities are consistently defined with respect to an inertial basis. For a total Lagrangian formulation, widely known rotation parametrizations used in rod finite element formulations are rotation vectors [undefl, undefag], quaternions [undefaa, undefaaq, undefaf, undefaan] and the 9-parameter method [undefh, undefaab, undefae]. Each parametrization poses additional requirement, such as a complementary update for rotation vectors, a constraint for the length of the quaternion and orthonormality constraints for the 9-parameter method.
@@
Furthermore, an additional challenge arises in the requirement for an objective interpolation of the kinematics. [undefp] discovered, that the direct interpolation of the nodal (total) rotation vectors leads to a non-objective formulation. Interpolating relative rotation vectors that describe the rotation relative to the first node solves this issue but introduces the additional computational step of evaluating these relative rotation vectors [[, see]]Jelenic1999.
@@
Another key challenge of rod finite element formulations is locking. As discussed in [undefd], shear and membrane locking in rods can occur if Kirchhoff (shear-rigidity) and inextensibility constraints follow in the limit case of a parameter tending to zero. This appears for instance for very slender rods if the stiffness parameters are computed in the sense of Saint–Venant by using the material’s Young’s and shear moduli, respectively, together with the cross-section geometry. Finite elements that are prone to locking cannot fulfill these constraints exactly over the entire element and introduce parasitic dilatation and shear strains. A common approach to mitigate this issue is reduced integration, which decreases deformation errors but fluctuations in resultant contact forces and moments are still present [[, see for instance]]Meier2015. A reinterpolation of the resultant contact forces and moments is possible to decrease the fluctuations. Other strategies to completely eliminate locking are B-bar methods [[, see for instance]]Greco2017, mixed formulations and three-field formulations based on the Hellinger–Reissner or the Hu–Washizu principle, respectively. Mixed and three-field formulations can be found in plate and shell finite element formulations [[, see]]Atluri1984,Betsch2016, Kirchhoff–Love rods [[, see]]Meier2014,Greco2024, or Cosserat rods [[, see]]Kim1998,Santos2010,Santos2011. In a planar study of arches, [undefak] use a cubic interpolation of the kinematic fields and quadratic interpolation for the resultant stresses together with the Bubnov–Galerkin approach. In [undefaah, undefaag], a so-called hybrid-mixed finite element formulation was derived, resulting in different polynomial degrees for the interpolation of orientation, displacement, force and moment. Mixed formulations can also be found in isogeometric analysis of Cosserat rods, see [undefaao, undefas], where the resultant contact forces and moments are interpolated with components in the inertial fixed basis and the collocation method is used. Also in rod formulations with enhanced kinematics, where warping or extension of the cross-section is taken into account, mixed formulations are applied, see for example [undefm, undefn].
@@
Mixed and three-field methods do not only solve the problem of locking, they can also significantly improve computational performance. This has been noticed by various authors and with different applications [undefap, undefao, undefay, undefas, undefe]. From a purely numerical perspective for solving systems of nonlinear equations, [undef] introduced the so-called lifted Newton method as a strategy to improve the convergence of Newton’s method, when applied to nonlinear root finding problems. Motivated by the multiple shooting method, they highlight, that a better rate of convergence can be achieved, if intermediate variables are introduced as additional degrees of freedom, constrained to their forward computed value. In a mixed formulation, the independent resultant contact forces and moments take the role of these intermediate variables. Since the equivalence of the mixed and displacement-based formulations can be shown under certain circumstances, as demonstrated in [undefar, undefax], the lifted Newton method arises naturally from the mixed formulation.
@@
Another effective way to remove locking is to employ intrinsically locking-free formulations, such as the SE(3){SE}(3) interpolation which has piece-wise constant strains. However, the equations are highly nonlinear, involve trigonometric functions for interpolation, couple position and rotation interpolations, and have computationally expensive evaluations of the force residuals and their derivatives. Additionally, these approaches can impose restrictions on scalability, since a refinement with a higher polynomial degree is not straightforwardly possible. While the authors of this paper used the SE(3){SE}(3) interpolation in a total Lagrangian framework [[, see]]Harsch2023, the interpolation can be done in the SE(3){SE}(3) Lie group setting [[, see]]Sonneville2014, requiring specialized Lie group solvers. [undefaaf] also presented a total Lagrangian SE(3){SE}(3) formulation. In contrast to the work of Harsch, where a Petrov–Galerkin projection was applied on the weak form of the Cosserat rod’s equilibrium equations, Santana derived a so-called equilibirum-based formulation, where the resultant contact forces and moments are obtained from the strong form of the rod’s differential equations. Furthermore, strain-based formulations should be mentioned, as they also can represent vanishing strains. They are parametrized directly by the strain measures instead of positions and orientations. A well known approach is the piece-wise constant strain formulation [[, see for instance]]Renda2018, which is widely used in the field of soft robotics. However, a major drawback of strain-based formulations is that positions and orientations can only be obtained by recursive algorithms, leading to asymmetric formulations.
@@
Beyond these numerical challenges, there is also the question of if and how to incorporate constrained rod theories within the general Cosserat rod framework. One approach, as explored in [undefae], introduces explicit constraints to enforce the Kirchhoff–Love assumptions. Of course, strain-based rod formulations can fulfill the constraints without any further effort, but the previously mentioned drawback remains. The common alternative is to intrinsically satisfy the Kirchhoff–Love assumption by advanced kinematic interpolation itself. Several formulations follow this path, yet they face similar challenges as Cosserat rods do, regarding objectivity, locking, and equation complexity. Moreover, intrinsically Kirchhoff–Love rods are often formulated within an updated Lagrangian framework, inheriting the associated numerical difficulties [[, see for instance]]Meier2016.
@@
The remainder of this paper is organized as follows. In Section˜2, the Cosserat rod theory is briefly recapitulated. It includes the spatially continuous formulations of the internal virtual work and the external virtual work. Two formulations for the internal virtual work are introduced. The first formulation uses only kinematics quantities, while the second is derived using the variational Hellinger–Reissner principle resulting in a formulation that incorporates kinematics and independent resultant contact forces and moments. In Section˜3, the Petrov–Galerkin rod finite element formulation of the Cosserat rod is derived. Therein, the ansatz functions for kinematics, the resultant contact forces and moments, and the corresponding test functions are discretized. The final rod finite element formulation is then obtained by inserting the discretized fields into the continuous formulation from Section˜2, leading to the equations of the generalized force equilibrium. Section˜4 is dedicated to comparing the obtained rod finite element formulations. The comparison includes the displacement-based SE(3){SE}(3) interpolation outlined in [undefag], and the proposed mixed quaternion formulation of this paper with different polynomial degrees. For completeness, we also consider a mixed SE(3){SE}(3) interpolation and displacement-based quaternion interpolations. Further, we investigate the influence of different spatial integration schemes. Well-studied benchmark experiments are conducted to show the validity and correctness of the proposed formulations. Finally, Section˜5 presents the concluding remarks. In Appendix˜A, the boundary value problem of the Cosserat rod is formulated. Appendix˜B gives further details on the finite element formulation. The detailed derivation of the tangent operator to the quaternion function for non-unit quaternions is given in Appendix˜C. Information on the numerical convergence analyses are given in Appendix˜D.
@@ 2 Cosserat rod theory
As in [undefaai, Equation 2.10], we introduce the diagonal elasticity matrices ${\bm{C}}_{{\bm{\gamma}}}=\operatorname{\operatorname{diag}}(k_{\mathrm{e}},k_{\mathrm{s}_{y}},k_{\mathrm{s}_{z}})$ and ${\bm{C}}_{{\bm{\kappa}}}=\operatorname{\operatorname{diag}}(k_{\mathrm{t}},k_{\mathrm{b}_{y}},k_{\mathrm{b}_{z}})$ with constant coefficients. In the following, the simple quadratic strain energy density
@@
Note that even in the inextensible case, where no strain energy density $W$ with finite stiffness coefficients is available, the complementary strain energy density ${W}^{\star}$ still has finite coefficients. This can be seen as the limit of stiffness components approaching infinity, which results in zeros on the diagonals of ${\bm{C}}_{{\bm{\gamma}}}^{-1}$ and ${\bm{C}}_{{\bm{\kappa}}}^{-1}$. Shear-rigidity can therefore easily be introduced by using ${\bm{C}}_{{\bm{\gamma}}}^{-1}=\operatorname{\operatorname{diag}}(k_{\mathrm{e}}^{-1},0,0)$. Additionally, adding the inextensibility constraint leads to ${\bm{C}}_{{\bm{\gamma}}}^{-1}=\mathbf{0}_{3\times 3}$. The same results were obtained by [undefv], where still the internal virtual work from ˜4 with the constitutive equations ˜3 was used for the unconstrained deformations.
@@ 3.1 Quaternion interpolation
Its derivation is shown in Appendix˜C. This kinematic interpolation was already proposed by [undefaac], who also provided a proof of objectivity. In contrast to the $\mathbb{R}^{12}$ interpolation introduced by [undefh] and [undefaab], the quaternion map in ˜13 ensures that ${\bm{A}}_{IB}$ is always an orthonormal transformation matrix.
@@ 3.3 Equations of static equilibrium
are obtained. Note that the number of equations equals in both formulations the number of unknowns, since ${\bm{q}}\in\mathbb{R}^{7N}$ and ${\bm{\lambda}}_{c}\in\mathbb{R}^{6pn_{\mathrm{el}}}$ with the relation $N=pn_{\mathrm{el}}+1$. Following [undefz], prescribed boundary conditions can be incorporated into the principle of virtual work using perfect bilateral constraints. In both formulations, the nonlinear equilibrium equations can be represented in the form ${\bm{f}}({\bm{x}})=\mathbf{0}_{n}$, where ${\bm{x}}\in\mathbb{R}^{n}$ contains all unknowns, i.e., the generalized coordinates, possibly Lagrange multipliers for bilateral constraints and in the mixed formulation also the generalized forces. Such a nonlinear equation is solved by any root-finding algorithm, e.g., Newton–Raphson, Riks. Note that a system of linear equations with a non-symmetric iteration matrix $\partial{\bm{f}}/\partial{\bm{x}}\in\mathbb{R}^{n\times n}$ must be solved in each iteration. In this work, the static equilibrium will be computed using Newton’s method. As termination criteria we require the Euclidean norm $\sqrt{{\bm{f}}({\bm{x}})^{\!\mathrm{T}}{\bm{f}}({\bm{x}})}$ to be smaller than $\varepsilon\sqrt{n}$, where $\varepsilon$ is the tolerance.
Although we derived only the equations for static equilibrium in this work, for dynamics the equations of motion can also be derived. Since the same generalized coordinates are chosen as in [undefaf] and also the interpolation of the variations is identical, the inertial virtual work and the kinematic differential equation can be derived in the same way, leading to equations of motion.